# Approaches to the Inverse Fourier Transformation with Limited and Discrete Data

### v1

@@ I Introduction
The determination of partonic structure from first principles remains one of the central challenges in quantum chromodynamics (QCD). Lattice QCD provides a systematically improvable approach to nonperturbative structures. However, the Euclidean formulation of lattice prevents a direct computation of correlation functions defined on the light cone, such as parton distribution functions (PDFs), light-cone distribution amplitudes (LCDAs), generalized parton distributions (GPDs), and transverse-momentum dependent distributions (TMDs). The traditional operator product expansion (OPE) techniques can only access a limited number of Mellin moments of structure functions. Recent years, the large-momentum effective theory (LaMET) Ji:2013dva; Ji:2014gla has opened a promising avenue by relating Euclidean “quasi-distributions” computed in large momenta boosted hadrons to their corresponding light-cone counterparts through factorization and perturbative matching Zhang:2017bzy; Zhang:2020gaj; Holligan:2023rex; Hua:2020gnw; LatticeParton:2022zqc; Baker:2024zcd; Cloet:2024vbv; Xiong:2013bka; Alexandrou:2016eyt; Chen:2017mie; Zhang:2017zfe; Xu:2018mpf; Liu:2018hxv; Wang:2019msf; Zhang:2019qiq; Chen:2020ody; Ji:2020brr; Ji:2020ect; LatticeParton:2020uhz; Lin:2020rxa; Bhattacharya:2021moj; Gao:2021hxl; LatticePartonLPC:2021gpi; Li:2021wvl; Deng:2022gzi; Gao:2022iex; Gao:2022uhg; LatticeParton:2022xsd; LatticePartonCollaborationLPC:2022myp; LatticePartonLPC:2022eev; Zhang:2022xuw; Zhu:2022bja; Deng:2023csv; Ji:2023pba; LatticeParton:2023xdl; Zhao:2023ptv; Liu:2023onm; Avkhadiev:2024mgd; Good:2024iur; Han:2024cht; Han:2024min; Holligan:2024umc; Holligan:2024wpv; Ji:2024hit; Wang:2024wwa; LatticeParton:2024zko; Zhang:2024omt; Bollweg:2025iol; Wang:2025uap; Ji:2025mvk. A closely related framework, the pseudo-distribution method Radyushkin:2017cyf; Orginos:2017kos, instead makes use of coordinate-space correlation functions evaluated at multiple hadron momenta, and reconstructs the light-cone distribution by matching and fitting the model functions in momentum spaceRadyushkin:2017lvu; Zhang:2018ggy; Karpie:2018zaz; Joo:2019jct; Joo:2019bzr; HadStruc:2021qdf; Bhat:2022zrw; Kovner:2024pwl; Bhattacharya:2024qpp; HadStruc:2024rix; NieMiera:2025inn.
@@
An essential numerical step in both the LaMET and pseudo-distribution approaches is the reconstruction of momentum-fraction distributions from nonlocal matrix elements computed in Euclidean lattice QCD. This problem is mathematically equivalent to performing a limited inverse discrete Fourier transform (L-IDFT), since the data are available only at finite and discrete non-local separations, and are inevitably contaminated by statistical noise. Recently, several works have addressed the problem of L-IDFT from different perspectives Chen:2025cxr; Xiong:2025obq; Dutrieux:2025jed; Medrano:2025cmg. In the hybrid scheme Ji:2020brr; Chen:2025cxr, it was shown that if lattice data can be obtained at sufficiently large nonlocal separations, one may employ physically motivated extrapolations to effectively avoid the inverse problem. In the recent study Xiong:2025obq for ill-posedness in limited discrete fourier inversion and its regularization, we provided a rigorous mathematical proof that the L-IDFT satisfies the Hadamard criteria of existence and uniqueness, but fails the stability condition: small perturbations in the input can lead to exponentially amplified deviations in the reconstructed distribution. This formulates the L-IDFT as a genuine inverse problem. We further demonstrated that it falls into the class of moderately tractable ill-posed problems, implying that established methods from the theory of inverse problems can be successfully applied to stabilize the reconstruction.
@@
This work presents a systematic comparison of several representative approaches for solving the limited inverse discrete Fourier transform problem. We develop a unified testing strategy to evaluate their performance, employing both controlled toy models and real lattice data relevant to LaMET studies of light cone distribution amplitudes (LCDAs). Our results show that, in most cases, the Tikhonov regularization, the Bayesian approach with GRW, and the feedforward artificial neural network can successfully reconstruct the quasi-distributions in momentum space. Among them, the Bayesian and ANN methods exhibit superior stability and accuracy. These investigations thus provide valuable methodological insights for future lattice QCD extractions of parton physics. The paper is organized as follows: In Sec. II, we provide a brief overview of the formulation of ill-posed inverse problems and introduce four approaches for addressing the limited inverse discrete Fourier transform: Tikhonov regularization, the Backus–Gilbert method, the Bayesian approach with GRW, and the feedforward artificial neural networks. In Sec. III, we present numerical tests of these four methods using two distinct toy models and one set of real lattice data. Finally, Sec. IV provides a summary and concluding remarks of the present study.
@@ II.1 Definition of Ill-posed Inverse Problems
II.2 Limited Discrete Fourier Inversion Problem
@@ III.3 Real data from lattice QCD simulation
To examine the robustness limits of the inverse-problem approach under varying data quality, we systematically investigate the breakdown point of the reconstruction as the $\lambda$-range is truncated. The detailed results and analysis are presented in Appendix A.
@@ IV Summary and outlook
Therefore, our work suggest that researchers—whether working within the LaMET framework or the pseudo-distribution method can treat the limited inverse discrete Fourier transform (L-IDFT) as a well-defined inverse problem that can be properly addressed. However, when the data suffer from excessively short $\lambda$ truncation or insufficient precision, the inverse problem becomes ill-conditioned to the extent that meaningful reconstruction is no longer feasible. In such cases, improving the precision of the lattice data itself is a prerequisite rather than refining the inversion method. In contrast, given data with sufficiently large $\lambda$ coverage and adequate precision, one should select an appropriate reconstruction strategy based on the characteristics of the data. A more comprehensive approach would involve employing multiple methods simultaneously to provide a robust estimate of the systematic uncertainties.

### v2

@@ I Introduction
The determination of partonic structure from first principles remains one of the central challenges in quantum chromodynamics (QCD). Lattice QCD provides a systematically improvable approach to nonperturbative structures. However, the Euclidean formulation of lattice prevents a direct computation of correlation functions defined on the light cone, such as parton distribution functions (PDFs), light-cone distribution amplitudes (LCDAs), generalized parton distributions (GPDs), and transverse-momentum dependent distributions (TMDs). The traditional operator product expansion (OPE) techniques can only access a limited number of Mellin moments of structure functions. Recent years, the large-momentum effective theory (LaMET) Ji:2013dva; Ji:2014gla has opened a promising avenue by relating Euclidean “quasi-distributions” computed in large momenta boosted hadrons to their corresponding light-cone counterparts through factorization and perturbative matching Zhang:2017bzy; Zhang:2020gaj; Holligan:2023rex; Hua:2020gnw; LatticeParton:2022zqc; Baker:2024zcd; Cloet:2024vbv; Xiong:2013bka; Alexandrou:2016eyt; Chen:2017mie; Zhang:2017zfe; Xu:2018mpf; Liu:2018hxv; Wang:2019msf; Zhang:2019qiq; Chen:2020ody; Ji:2020brr; Ji:2020ect; LatticeParton:2020uhz; Lin:2020rxa; Bhattacharya:2021moj; Gao:2021hxl; LatticePartonLPC:2021gpi; Li:2021wvl; Deng:2022gzi; Gao:2022iex; Gao:2022uhg; LatticeParton:2022xsd; LatticePartonCollaborationLPC:2022myp; LatticePartonLPC:2022eev; Zhang:2022xuw; Zhu:2022bja; Deng:2023csv; Ji:2023pba; LatticeParton:2023xdl; Zhao:2023ptv; Liu:2023onm; Avkhadiev:2024mgd; Good:2024iur; Han:2024cht; Han:2024min; Holligan:2024umc; Holligan:2024wpv; Ji:2024hit; Wang:2024wwa; LatticeParton:2024zko; Zhang:2024omt; Bollweg:2025iol; Wang:2025uap; Ji:2025mvk. A closely related framework, the pseudo-distribution method Radyushkin:2017cyf; Orginos:2017kos, instead makes use of coordinate-space correlation functions evaluated at multiple hadron momenta, and reconstructs the light-cone distribution by matching and fitting the model functions in momentum spaceRadyushkin:2017lvu; Zhang:2018ggy; Karpie:2018zaz; Joo:2019jct; Joo:2019bzr; HadStruc:2021qdf; Karpie:2021pap; Bhat:2022zrw; Kovner:2024pwl; Bhattacharya:2024qpp; HadStruc:2024rix; NieMiera:2025inn.
@@
An essential numerical step in both the LaMET and pseudo-distribution approaches is the reconstruction of momentum-fraction distributions from nonlocal matrix elements computed in Euclidean lattice QCD. This problem is mathematically equivalent to performing a limited inverse discrete Fourier transform (L-IDFT), since the data are available only at finite and discrete non-local separations, and are inevitably contaminated by statistical noise. Recently, several works have addressed the problem of L-IDFT from different perspectives Karpie:2019eiq; Dutrieux:2024rem; Chen:2025cxr; Xiong:2025obq; Dutrieux:2025jed; Medrano:2025cmg. In the hybrid scheme Ji:2020brr; Chen:2025cxr, it was shown that if lattice data can be obtained at sufficiently large nonlocal separations, one may employ physically motivated extrapolations to effectively avoid the inverse problem. In the recent study Xiong:2025obq for ill-posedness in limited discrete fourier inversion and its regularization, we provided a rigorous mathematical proof that the L-IDFT satisfies the Hadamard criteria of existence and uniqueness, but fails the stability condition: small perturbations in the input can lead to exponentially amplified deviations in the reconstructed distribution. This formulates the L-IDFT as a genuine inverse problem. We further demonstrated that it falls into the class of moderately tractable ill-posed problems, implying that established methods from the theory of inverse problems can be successfully applied to stabilize the reconstruction.
@@
This work presents a systematic comparison of several representative approaches for solving the limited inverse discrete Fourier transform problem. We develop a unified testing strategy to evaluate their performance, employing both controlled toy models and real lattice data relevant to LaMET studies of light cone distribution amplitudes (LCDAs). Our results show that, in most cases, the Tikhonov regularization, the Bayesian approach with GRW, and the feedforward artificial neural network can successfully reconstruct the quasi-distributions in momentum space. Among them, the Bayesian and ANN methods exhibit superior stability and accuracy. Our study and comparisons further demonstrate that the inverse Fourier transform problem, unlike those encountered in spectral analyses or in reconstructing distributions from OPE moments, is a moderately tractable ill-posed problem. When lattice data of sufficient quality are available at large nonlocal separations, this problem can indeed be effectively addressed. These investigations thus provide valuable methodological insights for future lattice QCD extractions of parton physics. The paper is organized as follows: In Sec. II, we provide a brief overview of the formulation of ill-posed inverse problems and introduce four approaches for addressing the limited inverse discrete Fourier transform: Tikhonov regularization, the Backus–Gilbert method, the Bayesian approach with GRW, and the feedforward artificial neural networks. In Sec. III, we present numerical tests of these four methods using two distinct toy models and one set of real lattice data. Finally, Sec. IV provides a summary and concluding remarks of the present study.
@@ II.1 Definition of Ill-posed Inverse Problems
II.2 Limited Inverse Discrete Fourier Transform
@@ III.3 Real data from lattice QCD simulation
To examine the robustness limits of the inverse-problem approach under varying data quality, we systematically investigate the breakdown point of the reconstruction as the $\lambda$-range is truncated. The detailed results and analysis are presented in Appendix. A.
@@ IV Summary and outlook
Therefore, our work suggest that researchers, whether working within the LaMET framework or the pseudo-distribution method can treat the limited inverse discrete Fourier transform (L-IDFT) as a well-defined inverse problem that can be properly addressed. Unlike the inverse problems encountered in lattice spectral analyses or in reconstructing distributions from OPE moments, this is a moderately tractable ill-posed problem, which can be effectively managed when the data quality at large nonlocal separations is sufficiently good. However, when the data suffer from excessively short $\lambda$ truncation or insufficient precision, the inverse FT becomes ill-conditioned to the extent that meaningful reconstruction is no longer feasible. In such cases, improving the precision of the lattice data itself is a prerequisite rather than refining the inversion method. In contrast, given data with sufficiently large $\lambda$ coverage and adequate precision, one can select an appropriate reconstruction strategy based on the characteristics of the data. A more comprehensive approach would involve employing multiple methods simultaneously to provide a robust estimate of the systematic uncertainties.
V Acknowledgments
We thank Andreas Schäfer, Yong Zhao and Xiangdong Ji for valuable discussions and constructive suggestions on this work. We also thank Ting Wei and Xiong-Bin Yan for valuable discussions in mathematics. We thank the LPC collaboration for providing us the lattice data of pion distribution amplitudes LatticeParton:2022zqc. This work is supported in part by Natural Science Foundation of China under grant No. 12205106, 12375069 and 12335003. M. H. C. is supported by the National Science Centre (Poland) grant OPUS No. 2021/43/B/ST2/00497. Q.A.Z is supported by the Fundamental Research Funds for the Central Universities. A.S.X is supported the Fundamental Research Funds for the Central Universities under No. lzujbky-2023-stlt01.